# Linear and nonlinear instability of vortex columns

### v2

@@ Abstract.
We consider vortex column solutions $v=V(r)e_{\theta}+W(r)e_{z}$ to the $3$D Euler equations. We give a mathematically rigorous construction of the countable family of unstable modes discovered by Liebovich and Stewartson (J. Fluid Mech. 126, 1983) via formal asymptotic analysis. The unstable modes exhibit $O(1)$ growth rates and concentrate on a ring $r=r_{0}$ asymptotically as the azimuthal and axial wavenumbers $n,\alpha\to\infty$ with a fixed ratio. We construct these so-called ring modes with an inner-outer gluing procedure. Finally, we prove that each linear instability implies nonlinear instability for vortex columns. In particular, our analysis yields nonlinear instability for the Batchelor trailing line vortex $V(r):=\frac{q}{r}(1-\mathrm{e}^{-r^{2}})$ and $W(r):=\mathrm{e}^{-r^{2}}$ when $0<q\ll 1$.
@@ 1. Introduction
Vortex columns are steady solutions to the $3$D Euler equations
$\begin{split}\partial_{t}v+(v\cdot\nabla)v+\nabla p&=0\\ \operatorname{div}\,v&=0\end{split}$ (1.1)
@@
Vortices are among the simplest possible flows, and it is a classical goal of hydrodynamic stability theory to understand under what conditions these flows are stable. For vortices without axial flow, the most well known conditions are due to Rayleigh [45, 46] (see also Synge [49]); with axial flow, analogous conditions were established by Howard and Gupta [29]. The stability problem for vortex columns is further motivated by vortex breakdown, a nonlinear phenomenon historically characterized by the development of a stagnation point on the axis and a region of reversed flow as the vortex transitions to a more complex unsteady flow downstream, see [32, 33, 41] and [51, p. 74]. It is natural to expect that instabilities of the underlying vortex play a role in triggering its breakdown.
@@ 1.1. The Rayleigh equation for vortex columns
This equation can be block diagonalized by applying the Fourier transform in the $z$ variable and Fourier series in the $\theta$ variable. It will therefore be enough to consider solutions
@@
with eigenvalue $\lambda=-i\omega$, where we abuse notation by reusing $u_{r}$, etc. Notice the notational convention of $-n$ in (1.8) which is meant to conform with [34]. We will refer to unstable modes as nontrivial solutions $u$ of the form (1.8) with $\mathrm{Re}\,\lambda>0$, i.e.
@@
where [^2] In the definition of $a$ in [34, (4.3b)], there is a typo ($qn$ instead of $q$), which is corrected in [35, (2.3)].
@@
$n^{2}\left(\omega\left(\frac{1}{r}+\beta^{2}r\right)+V\left(\frac{n}{r^{2}}+n\beta^{2}\right)-W\left(\frac{\alpha}{r}+\alpha\beta^{2}r\right)\right)$
@@
The above Rayleigh equation is more complicated than its analogue for shear flows and vortices (see [4, 37, 38, 52, 53]), but it is still possible to extract some stability conditions from it. Notably, there is a necessary and sufficient condition for stability with respect to a restricted class of perturbations. Namely, Howard and Gupta [29] demonstrated that a columnar vortex is spectrally stable to axisymmetric perturbations, that is, $n=0$, precisely when its Rayleigh function $\Phi$ is non-negative. A second interesting special case is when $W=0$, that is, when we consider 3D perturbations of a 2D vortex. When the 2D vortex satisfies the Rayleigh condition [46], namely, that its vorticity profile is decreasing, Gallay and Smets demonstrated spectral stability in the enstrophy class [24] and, in the energy class, that the growth rate of the linear semigroup is, at most, subexponential [23]. We refer to the review article [22] for further discussion. We discuss additional stability criteria for vortex columns (1.2), including an analogue of the semicircle theorem [6], in Section 1.2.1.
@@ 1.2. Linear instability
in which case $q$, defined in (1.10), is the constant $q=\max_{r}\Omega(r)$ representing the maximal pitch angle of the helical particle trajectories of the basic flow. We show in Section A.1 that for every sufficiently small $q>0$ there exists a unique choice of $r_{0},\beta>0$ satisfying (1.14), and we verify a necessary condition
$\frac{q}{2}<\beta<\frac{1}{q}.$ (1.15)
The two limiting cases $\beta\to(1/q)^{-}$ and $\beta\to(q/2)^{+}$, were studied by Capell and Stewartson [10], and Leibovich and Stewartson [35], respectively.
@@
$\phi(r):=\left(\frac{r^{3}}{1+\beta^{2}r^{2}}\right)^{\frac{1}{2}}u(r)\,.$ (1.16)
@@
We now summarize the key points of the asymptotics analysis of Leibovich and Stewartson [34]. They sought a sequence of unstable modes with growth rates $\operatorname{Im}\omega=O(1)$ as $n\to+\infty$ and the wavenumber ratio $\beta=\alpha/n$ is constant. For unstable modes to exist, it is necessary that the potential $k$, defined in (1.18), be somewhere negative. It is not obvious that this is possible, due to the presence of the “$1$” in the bracket in (1.18). It is plausible that the terms $a/(n\gamma)$ and $d/n^{2}$, containing $1/n$, will be negligible compared to $1+b/\gamma^{2}$, and we temporarily ignore them. (Recall (1.11) that $\gamma=n\Lambda-\omega$.) For $b/\gamma^{2}$ not to be negligible, it is necessary that $\omega=n\Lambda(r_{0})+o(n)$ for a value $r_{0}$. Subsequently, we suppose that
$\omega=\underbrace{n\Lambda(r_{0})+ib(r_{0})^{1/2}}_{=:\omega_{\rm app}}+o(1),$ (1.20)
@@
$k(r)=k_{0}+k_{2}(r-r_{0})^{2}+\text{\emph{ lower order terms}}\,,$ (1.21)
where the requirement $\Lambda^{\prime}(r_{0})=\gamma^{\prime}(r_{0})=0$ eliminates the $k_{1}$ term, and the lower order terms may depend on $r$ and $n$. A direct computation shows that
@@
for some $m\in{\mathbb{N}}$, in which case the exact solutions are constant multiples of the $m$th Weber function
$w_{m}(r):=c_{m}\exp\left(-\frac{1}{2}k_{2}^{1/2}(r-r_{0})^{2}\right)H_{m}\left(k_{2}^{1/4}(r-r_{0})\right)$ (1.25)
where $H_{m}(x):=(-1)^{m-1}\mathrm{e}^{x^{2}}(\frac{\mathrm{d}}{\mathrm{d}x})^{m-1}\mathrm{e}^{-x^{2}}$ denotes the $m$th Hermite polynomial and $c_{m}>0$ is a normalization factor ensuring $n^{3/8}\|w_{m}\|_{L^{2}}=1$. The branch of $k_{2}^{1/2}$ is chosen so that $\mathrm{Re}\,k_{2}^{1/2}>0$, i.e.,
@@
$\omega_{m}=n\Lambda_{0}+ib_{0}^{1/2}+\underbrace{(1-i)n^{-1/2}(2m-1)\left(\frac{b_{0}^{1/2}\Lambda^{\prime\prime}(r_{0})}{8p_{0}}\right)^{1/2}}_{=:\mu_{m}}+\widehat{\omega}_{m}\,,$ (1.27)
@@
The above choice sets the “inner length scale” as $O(n^{-3/4})$, as can be seen from the scaling factor $n^{3/4}$ hiding inside $k_{2}^{1/4}$ in (1.25). In conclusion, we expect the $m$-th family of unstable modes to look like the $m$-th Weber function on length scale $O(n^{-3/4})$ and be centered at $r=r_{0}$. The eigenvalue should satisfy the asymptotics (1.27). In [34], the authors present an impressive (formal) calculation which yields the eigenvalue up to and including the $O(n^{-3/2})$ term of $\omega$ for the “fundamental” mode $m=1$.
@@
$V\frac{d\Omega}{dr}\left[\frac{d\Omega}{dr}\frac{d\Gamma}{dr}+\left(\frac{dW}{dr}\right)^{2}\right]<0\text{ somewhere in the flow field }$ (1.28)
@@
The uniqueness assertion in Theorem 1.1 provides an asymptotic classification of sequences of unstable eigenfunctions $\phi_{k}$ and eigenvalues $-i\omega_{k}$ satisfying $|\omega_{k}-(n_{k}\Lambda_{0}+ib_{0}^{1/2})|\lesssim n_{k}^{-1/2}$ and $n_{k}\to+\infty$ as $k\to+\infty$.
@@ 1.2.1. Instability criteria
$\omega_{0}=n\Lambda(r_{0})/2\,,\qquad\omega_{\rm max}=\sqrt{n^{2}\Lambda(r_{0})^{2}/4+\max\widetilde{R}(r)^{2}}\,,$
@@
Figure 1. Unstable modes (2.27) and the Barston [6] semicircle theorem. As required by the semicircle theorem, $\sqrt{b_{0}}<\max_{r}\widetilde{R}$, i.e., all the modes obtained in Theorem 1.1 belong to the semicircle $B(\omega_{0},\omega_{\rm max})\cap\{\mathrm{Im}\,\omega>0\}$. In the case $q=1/4$, one has $\beta\approx 0.1589$, $r_{0}\approx 0.8252$ (see Appendix A.1), which then shows that $\sqrt{b_{0}}\approx 0.194<0.218\approx\max_{r}\widetilde{R}$.
We also point out the criterion of Howard and Gupta [29, (21)] that the flow (1.2) is stable if
@@
$-\left(1-\mathrm{e}^{-r^{2}}\right)(1+\beta q)\mathrm{e}^{-r^{2}}\frac{4q\beta}{r^{2}}-\frac{q^{2}}{r^{6}}\left(1-\mathrm{e}^{-r^{2}}\left(1+r^{2}+\frac{\beta}{q}r^{4}\right)\right)^{2}\geq 0\,,$ (1.32)
which is actually violated for all $r>0$; the second term on the left-hand side equals $-\frac{1}{4}\Lambda^{\prime}(r)^{2}$, which vanishes at $r_{0}$, the center of concentration of the ring modes.
@@
see also the discussion in [34, Section 5]. Furthermore, [29, (23)] provides an equivalent of the Rayleigh criterion for instability,
@@
Considering the scenario of neutral limiting modes (i.e., when $\mathrm{Im}\,\omega>0$ is obtained as a perturbation of $\mathrm{Im}\,\omega=0$), we have that, for the limiting mode, $\gamma$ is a real function, and one can deduce from (1.33) that, if $\left(\beta rV+W\right)^{\prime}$ vanishes, then
@@
However, this raises an interesting open problem of existence of the upper limiting modes as $\beta\to(1/q)^{-}$, that is, when the relation between $\alpha$ and $n$ is not constant. In that case, the analysis of this paper breaks down. Stewartson and Capell [10] provide some evidence of existence of such modes. However, the question of a rigorous verification and establishing a connection with the well-understood theory of instabilities bifurcating from neutral limiting modes (see [37] in the shear flow case) remains an interesting open problem. See [34] for additional information.
@@ 1.3. Nonlinear instability
In Theorem 1.2, we assume modal linear instability, that is, existence of an unstable solution to the linearization of the $3$D Euler equations (1.1) in the form (1.8), and prove that such linear instability gives rise to nonlinear instability.
@@
We fix $Q\in(1,2)$ and consider the linearization of the $3$D Euler equations (1.1),
@@
$-{\mathcal{L}}\omega:=[{\overline{u}},\omega]-[u,{\overline{\omega}}]={\overline{u}}\cdot\nabla\omega-\omega\cdot\nabla{\overline{u}}-{\overline{\omega}}\cdot\nabla u+u\cdot\nabla{\overline{\omega}}.$ (1.37)
Here we use the Lie bracket notation $[v,w]:=v\cdot\nabla w-w\cdot\nabla v$ for vector fields, and we also denote by $u$ the velocity field generated by $\omega$ via the Biot-Savart law. The definition of the Biot-Savart law on $\mathbb{R}^{2}\times\mathbb{T}$ is a subtle matter, since the domain $\mathbb{R}^{2}\times\mathbb{T}$ allows, in particular, two dimensional flows (i.e. constant in $z$). In particular, according to the two-dimensional Biot-Savart law ${\rm BS}_{2}:L^{2}\to\dot{H}^{1}$, the velocity fields are only defined up to constants, unless further conditions are imposed, e.g., more decay at spatial infinity; this is why we work in the space $L^{Q}$.
Let
[…]
which decomposes into invariant subspaces
$X_{n,\alpha}:=\{u\in L^{Q}_{\sigma}\colon u=v(r)\mathrm{e}^{in\theta+i\alpha z}\text{ for some }n,\alpha\in\mathbb{Z}\text{ and }v\in L^{Q}(\mathbb{R}^{+},r\mathrm{d}r)\}\,.$ (1.38)
@@
The existence of infinitely many such $\eta$s and $\lambda$s follows from Theorem 1.1 for any vortex column (1.2) for which the sufficient condition (1.14) holds, see Corollary 3.4. The size of the torus does not matter; after rescaling, the period can be assumed to be one.
@@
Suppose that a vortex column (1.34) satisfying the decay assumption (1.35) is linearly unstable in the sense of (1.40). Then the vortex column is nonlinearly unstable in the $L^{\infty}(\mathbb{R}^{2}\times\mathbb{T})$ topology of velocity in the following sense: For every $N\geq 2$ there exists $\delta>0$ such that, for any $\varepsilon>0$, there exist $T_{\varepsilon}>0$ and a strong solution $u$ of the $3$D Euler equations (1.1) on $\mathbb{R}^{2}\times\mathbb{T}\times[0,T_{\varepsilon}]$ such that
@@
In Section 2, we construct the unstable eigenfunctions advertised in Theorem 1.1. In Section 3, we prove the nonlinear instability in Theorem 1.2.
@@
keeping in mind that the existence proof of Theorem 1.1 is assumed for a fixed $m$.
@@
The ansatz involves some redundancy of the ‘inner solution’ $\phi_{\rm in}$ and ‘outer solution’ $\phi_{\rm out}$ on the overlap region; this is typical of gluing procedures. Then $\phi^{\prime\prime}-k\phi=0$ if
@@
provided that $|\widetilde{\omega}|+|r-r_{0}|=O(n^{-1/2})$, see (A.3) in the appendix for the precise statement and proof. We rescale the potential $k$ according to
$\begin{split}K(\xi)&:=n^{-3/2}k(r),\quad K_{0}:=n^{-3/2}k_{0},\quad K_{2}:=n^{-3/2}k_{2}\\ K_{\rm err}(\xi)&:=n^{-3/2}k_{\rm err}(r)=K(\xi)-(K_{0}+K_{2}\xi^{2})\,.\end{split}$ (2.8)
@@
${\Phi}_{\rm in}^{\prime\prime}-(K_{0}+K_{2}\xi^{2}){\Phi}_{\rm in}=K_{\rm err}{\Phi}_{\rm in}-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\text{ on }{\rm supp}\,X_{\rm in}\,.$ (2.12)
@@
are precisely constant multiples of the the Weber functions $W_{m}=R^{-1}w_{m}$, recall (1.25), which implies in particular that $\|W_{m}\|_{L^{2}}=1$. As mentioned in the beginning of the section, we will drop the index $m$. We make the ansatz
@@
The operator $\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}-(K_{0}+K_{2}\xi^{2})$ has a one-dimensional kernel spanned by $W$, and its range is $({\rm span}\,\overline{W})^{\perp}$ (see Lemma 2.2).555 Multiplying (2.16) equation by $W$ yields that $\int WG=0$, which suggest that this is a solvability condition for $\Psi$. The equation (2.16) is therefore a suitable candidate for the Lyapunov-Schmidt reduction. We thus define the projection $Q\colon L^{2}\to L^{2}$ onto the range:
$QF:=F-\left(\int FW\right)\overline{W}\,,$ (2.17)
and we rewrite (2.15) as the projected equation

### v3

@@ Abstract.
We consider vortex column solutions $v=V(r)e_{\theta}+W(r)e_{z}$ to the 33D Euler equations. We give a mathematically rigorous construction of the countable family of unstable modes discovered by Liebovich and Stewartson (J. Fluid Mech. 126, 1983) via formal asymptotic analysis. The unstable modes exhibit $O(1)$ growth rates and concentrate on a ring $r=r_{0}$ asymptotically as the azimuthal and axial wavenumbers $n,\alpha\to\infty$ with a fixed ratio. We construct these so-called ring modes with an inner-outer gluing procedure. Finally, we prove that each linear instability implies nonlinear instability for vortex columns. In particular, our analysis yields nonlinear instability for the Batchelor trailing line vortex $V(r):=\frac{q}{r}(1-\mathrm{e}^{-r^{2}})$ and $W(r):=\mathrm{e}^{-r^{2}}$ when $0<q<\log 2/\sqrt{1-\log 2}\approx 1.251$.
@@ 1. Introduction
Vortex columns are steady solutions to the 33D Euler equations
$\begin{split}\partial_{t}v+(v\cdot\nabla)v+\nabla p&=0,\\ \operatorname{div}\,v&=0\end{split}$ (1.1)
@@
Vortices are among the simplest possible flows, and it is a classical goal of hydrodynamic stability theory to understand under what conditions these flows are stable. For vortices without axial flow, the most well known conditions are due to Rayleigh [47, 48] (see also Synge [51]); with axial flow, analogous conditions were established by Howard and Gupta [30]. The stability problem for vortex columns is further motivated by vortex breakdown, a nonlinear phenomenon historically characterized by the development of a stagnation point on the axis and a region of reversed flow as the vortex transitions to a more complex unsteady flow downstream, see [34, 35, 43] and [53, p. 74]. It is natural to expect that instabilities of the underlying vortex play a role in triggering its breakdown.
@@ 1.1. The Rayleigh equation for vortex columns
These can be block diagonalized by applying the Fourier transform in the $z$ variable and Fourier series in the $\theta$ variable. It will therefore be enough to consider solutions
@@
with eigenvalue $\lambda=-i\omega$, where we abuse notation by reusing $u_{r}$, etc. Namely, we seek solutions of the form $u\mathrm{e}^{-i\omega t}$ to the linearized Euler equations, which are equivalent to the system (1.7) for $u$ and $\lambda=-i\omega\in\mathbb{C}$. Notice the notational convention of n-n in (1.8) which is meant to conform with [36]. We will refer to unstable modes as nontrivial solutions $u$ of the form (1.8) with $\mathrm{Re}\,\lambda>0$, i.e.
@@
where [^2] In the definition of $a$ in [36, (4.3b)], there is a typo ($qn$ instead of $q$), which is corrected in [37, (2.3)].
@@
$n^{2}\left(\omega\left(\frac{1}{r}+\beta^{2}r\right)+\left(\frac{n}{r^{2}}+n\beta^{2}\right)V-\left(\frac{\alpha}{r}+\alpha\beta^{2}r\right)W\right)$
@@
The above Rayleigh equation (1.9) is more complicated than its analogue for shear flows and vortices (see [4, 33, 39, 40, 54, 55]), but it is still possible to extract some stability conditions from it. Notably, there is a necessary and sufficient condition for stability with respect to a restricted class of perturbations. Namely, Howard and Gupta [30] demonstrated that a columnar vortex is spectrally stable to axisymmetric perturbations, that is, $n=0$, precisely when its Rayleigh function $\Phi$ is non-negative. A second interesting special case is when $W=0$, that is, when we consider 3D perturbations of a 2D vortex. When the 2D vortex satisfies the Rayleigh condition [48], namely, that its vorticity profile is decreasing, Gallay and Smets demonstrated spectral stability in the enstrophy class [24] and, in the energy class, that the growth rate of the linear semigroup is, at most, subexponential [23]. We refer to the review article [22] for further discussion; in Section 1.2.1 we also discuss additional stability criteria for vortex columns (1.2), including an analogue [6] of Howard’s semicircle theorem [29].
@@ 1.2. Linear instability
in which case $q$, defined in (1.10), is the constant $q=\max_{r}\Omega(r)$ representing the maximal pitch angle of the helical particle trajectories of the basic flow. We show in Appendix A.1 that
$\frac{q}{2}<\beta<\frac{1}{q}$ (1.15)
is a necessary condition for Assumption A, and we show that there exist a unique choice of $r_{0},\beta>0$ satisfying (1.14) if
$q<\frac{\log 2}{\sqrt{1-\log 2}}\approx 1.251.$ (1.16)
Note that (1.15) implies that $q<2^{1/2}$, but, as shown in Appendix A.1, for $q$ close to $2^{1/2}$ there are no $r_{0},\beta$ satisfying (1.14). We also note that the two limiting cases $\beta\to(1/q)^{-}$ and $\beta\to(q/2)^{+}$, were studied by Capell and Stewartson [10] and Leibovich and Stewartson [37], respectively.
@@
$\phi(r):=\left(\frac{r^{3}}{1+\beta^{2}r^{2}}\right)^{\frac{1}{2}}u_{r}(r)\,.$ (1.17)
@@
We now summarize the key points of the asymptotics analysis of Leibovich and Stewartson [36]. They sought a sequence of unstable modes with growth rates $\operatorname{Im}\omega=O(1)$ as $n\to+\infty$ and the wavenumber ratio $\beta=\alpha/n$ is constant. For unstable modes to exist, it is necessary that the real part of the potential $k$, defined in (1.19), be somewhere negative. It is not obvious that this is possible, due to the presence of the “11” in the bracket in (1.19). It is plausible that the terms $a/(n\gamma)$ and $d/n^{2}$, containing $1/n$, will be negligible compared to $1+b/\gamma^{2}$, and we temporarily ignore them. (Recall (1.11) that $\gamma=n\Lambda-\omega$.) For $b/\gamma^{2}$ not to be negligible, it is necessary that $\omega=n\Lambda(r_{0})+o(n)$ as $n\to\infty$, for a value $r_{0}$. Subsequently, we suppose that
$\omega=\underbrace{n\Lambda(r_{0})+ib(r_{0})^{1/2}}_{=:\omega_{\rm app}}+o(1)\qquad\text{ as }n\to\infty,$ (1.21)
@@
$k(r)=k_{0}+k_{2}(r-r_{0})^{2}+\text{\emph{remainder}}\,$, (1.22)
where the the remainder may depend on $r$ and $n$, and the requirement $b^{\prime}(r_{0})=\Lambda^{\prime}(r_{0})=\gamma^{\prime}(r_{0})=0$ eliminates the $k_{1}$ term in the sense that the linear part of the expansion arises only from $p,a,d$ (recall (1.19)), which can be treated as part of the remainder. A direct computation shows that
@@
for some $m\in{\mathbb{N}}$, in which case the exact solutions are constant multiples of the $m$-th Weber function
$w_{m}(r):=c_{m}\exp\left(-\frac{1}{2}k_{2}^{1/2}(r-r_{0})^{2}\right)H_{m}\left(k_{2}^{1/4}(r-r_{0})\right),$ (1.26)
where $H_{m}(x):=(-1)^{m-1}\mathrm{e}^{x^{2}}(\frac{\mathrm{d}}{\mathrm{d}x})^{m-1}\mathrm{e}^{-x^{2}}$ denotes the $m$-th Hermite polynomial and $c_{m}>0$ is a normalization factor ensuring $n^{3/8}\|w_{m}\|_{L^{2}}=1$. Here the branch of $k_{2}^{1/2}$ is chosen so that $\mathrm{Re}\,k_{2}^{1/2}>0$, i.e.,
@@
$\omega_{m}=\underbrace{n\Lambda_{0}+ib_{0}^{1/2}}_{=\omega_{\rm app}}+\underbrace{(1-i)n^{-1/2}(2m-1)\left(\frac{b_{0}^{1/2}\Lambda^{\prime\prime}(r_{0})}{8p_{0}}\right)^{1/2}}_{=:\mu_{m}}+\widehat{\omega}_{m}\,$, (1.28)
@@
The above choice sets the “inner length scale” as $O(n^{-3/4})$, as can be seen from the scaling factor $n^{3/4}$ hiding inside $k_{2}^{1/4}$ in (1.26). In conclusion, we expect the $m$-th family of unstable modes to look like the $m$-th Weber function on length scale $|r-r_{0}|\leq O(n^{-3/4})$ and be centered at $r=r_{0}$. The eigenvalue $\omega$ (which is a part of the definition (1.19) of potential $k$) should satisfy the asymptotics (1.28). In [36], the authors present an impressive (formal) calculation which yields the eigenvalue up to and including the $O(n^{-3/2})$ term of $\omega$ for the “fundamental” mode $m=1$.
@@
$V\frac{\mathrm{d}\Omega}{\mathrm{d}r}\left[\frac{\mathrm{d}\Omega}{\mathrm{d}r}\frac{\mathrm{d}\Gamma}{\mathrm{d}r}+\left(\frac{\mathrm{d}W}{\mathrm{d}r}\right)^{2}\right]<0\text{ somewhere in the flow field }$ (1.29)
@@
The uniqueness assertion in Theorem 1.1 provides an asymptotic classification of sequences of unstable eigenfunctions $\phi_{k}$ and eigenvalues $i\omega_{k}$ satisfying $|\omega_{k}-(n_{k}\Lambda_{0}+ib_{0}^{1/2})|\lesssim n_{k}^{-1/2}$ and $n_{k}\to+\infty$ as $k\to+\infty$. Note that $C_{M}$ is such that $B(\omega_{\rm app},C_{M}n^{-1/2})$ contains the first M1M-1 modes, see Figure 1 below.
@@ 1.2.1. Instability criteria
$\omega_{0}=n\Lambda(r_{0})/2\,,\qquad\omega_{\rm max}=\sqrt{n^{2}\Lambda(r_{0})^{2}/4+\max_{r}\widetilde{R}(r)^{2}}\,$,
@@
Figure 1. Unstable modes (2.27) and the Barston [6] semicircle theorem. As required by the semicircle theorem, $\sqrt{b_{0}}<\max_{r}\widetilde{R}$, i.e., all the modes obtained in Theorem 1.1 belong to the semicircle $B(\omega_{0},\omega_{\rm max})\cap\{\mathrm{Im}\,\omega>0\}$. For example, in the case $q=1/4$, one has $\beta\approx 0.1589$, $r_{0}\approx 0.8252$ (see Appendix A.1), which then shows that $\sqrt{b_{0}}\approx 0.194<0.218\approx\max_{r}\widetilde{R}$.
We also point out the criterion of Howard and Gupta [30, (21)] that the flow (1.2) is stable if
@@
$\left(1-\mathrm{e}^{-r^{2}}\right)(1+\beta q)\mathrm{e}^{-r^{2}}\frac{4q\beta}{r^{2}}-\frac{q^{2}}{r^{6}}\left(1-\mathrm{e}^{-r^{2}}\left(1+r^{2}+\frac{\beta}{q}r^{4}\right)\right)^{2}\geq 0\,$, (1.33)
which is actually violated for all $r>0$; the second term on the left-hand side (including the minus sign) equals $\frac{1}{4}\Lambda^{\prime}(r)^{2}$, which vanishes at $r_{0}$, the center of concentration of the ring modes.
@@
see also the discussion in [36, Section 5]. Furthermore, [30, (23)] provides an analogue of the Rayleigh criterion for instability,
@@
Considering the scenario of neutral limiting modes (i.e., when $\mathrm{Im}\,\omega>0$ is obtained as a perturbation of $\mathrm{Im}\,\omega=0$), we have that, for the limiting mode, $\gamma$ is a real function (recall (1.11)), and one can deduce from (1.34) that, if $\left(\beta rV+W\right)^{\prime}$ vanishes, then
@@
However, this raises an interesting open problem of existence of the upper limiting modes as $\beta\to(1/q)^{-}$, that is, when the relation between $\alpha$ and $n$ is not constant. In that case, the analysis of this paper breaks down (as the assumption $b(r_{0})>0$ in (1.14) fails). Stewartson and Capell [10] provide some evidence of existence of such modes. However, the question of a rigorous verification and establishing a connection with the well-understood theory of instabilities bifurcating from neutral limiting modes (see [33, 39] in the shear flow case) remains an interesting open problem. See [36] for additional information.
@@ 1.3. Nonlinear instability
In our second result (Theorem 1.2 below) we deduce nonlinear instability from the modal instability. Namely, we assume existence of an unstable solution to the linearization of the 33D Euler equations (1.1) in the form (1.8), and prove that such linear instability gives rise to nonlinear instability.
@@
We focus on the domain $\mathbb{R}^{2}\times\mathbb{T}$ for convenience, due to the periodicity in $z$ of the modes of the form (1.8). After rescaling, we may take $\mathbb{T}$ to be the flat torus of unit length.
@@
We fix $Q\in(1,2)$ and consider the linearization of the 33D Euler equations (1.1),
@@
${\mathcal{L}}\omega:=[{\overline{u}},\omega]-[u,{\overline{\omega}}]={\overline{u}}\cdot\nabla\omega-\omega\cdot\nabla{\overline{u}}-{\overline{\omega}}\cdot\nabla u+u\cdot\nabla{\overline{\omega}}.$ (1.38)
Here we use the Lie bracket notation $v,w]:=v\cdot\nabla w-w\cdot\nabla v$ for vector fields, and we also denote by $u$ the velocity field generated by $\omega$ via the Biot-Savart law. The definition of the Biot-Savart law on $\mathbb{R}^{2}\times\mathbb{T}$ is a subtle matter, since the domain $\mathbb{R}^{2}\times\mathbb{T}$ allows, in particular, two dimensional flows (i.e. constant in $z$). In particular, according to the two-dimensional Biot-Savart law ${\rm BS}_{2}:L^{2}\to\dot{H}^{1}$, the velocity fields are only defined up to constants, unless further conditions are imposed, e.g., more decay at spatial infinity; this is why we work in the space $L^{Q}$. For any $p\geq 1$, we set
$L^{p}_{\sigma}:=\{f\in L^{p}(\mathbb{R}^{2}\times\mathbb{T};\mathbb{R}^{3})\colon f\text{ is weakly divergence-free}\}\,.$ (1.39)
We will consider a decomposition of $L^{Q}_{\sigma}$ into invariant subspaces
$X_{n,\alpha}:=\{u\in L^{Q}_{\sigma}\colon u=v(r)\mathrm{e}^{in\theta+i\alpha z}\text{ for some }n,\alpha\in\mathbb{Z}\text{ and }v\in L^{Q}((\mathbb{R}^{+},r\mathrm{d}r);\mathbb{R}^{3})\}\,.$ (1.40)
@@
[…]
@@
Suppose that a vortex column (1.35) satisfying the decay assumption (1.36) is linearly unstable in the sense of (1.42). Then the vortex column is nonlinearly unstable in the $L^{\infty}(\mathbb{R}^{2}\times\mathbb{T})$ topology of velocity in the following sense: For every $N\geq 2$ there exists $\delta>0$ such that, for any $\varepsilon>0$, there exist $T_{\varepsilon}>0$ and a strong solution $u$ of the 33D Euler equations (1.1) on $\mathbb{R}^{2}\times\mathbb{T}\times[0,T_{\varepsilon}]$ such that
@@
In Section 2, we construct the unstable eigenfunctions advertised in Theorem 1.1. In Section 3, we prove the nonlinear instability of Theorem 1.2.
@@
keeping in mind that the existence proof of Theorem 1.1 is carried out for a fixed $m$.
@@
The ansatz involves some redundancy of the ‘inner solution’ $\phi_{\rm in}$ and ‘outer solution’ $\phi_{\rm out}$ on the overlap region; this is typical of gluing procedures. The Rayleigh equation $\phi^{\prime\prime}-k\phi=0$ (recall (1.18)) holds if
@@
provided that $|\mu_{m}|+|r-r_{0}|=O(n^{-1/2})$, see (A.3) in the appendix for the precise statement and proof. We rescale the potential $k$ according to
$\begin{split}K(\xi)&:=n^{-3/2}k(r),\quad K_{0}:=n^{-3/2}k_{0},\quad K_{2}:=n^{-3}k_{2}\\ K_{\rm err}(\xi)&:=n^{-3/2}k_{\rm err}(r)=K(\xi)-(K_{0}+K_{2}\xi^{2})\,.\end{split}$ (2.8)
@@
${\Phi}_{\rm in}^{\prime\prime}-(K_{0}+K_{2}\xi^{2}){\Phi}_{\rm in}=K_{\rm err}{\Phi}_{\rm in}-R\phi_{\rm out}X_{\rm out}^{\prime\prime}-2n^{-3/4}R\phi_{\rm out}^{\prime}X_{\rm out}^{\prime}\text{ on }{\rm supp}\,X_{\rm in}\,.$ (2.12)
@@
are precisely constant multiples of the the Weber functions $W_{m}=Rw_{m}$, recall (1.26), which implies in particular that $\|W_{m}\|_{L^{2}}=1$. As mentioned in the beginning of the section, we will drop the index $m$. We make the ansatz
@@
The operator $\frac{\mathrm{d}^{2}}{\mathrm{d}\xi^{2}}-(K_{0}+K_{2}\xi^{2})$ has a one-dimensional kernel spanned by $W$, and its range is $({\rm span}\,W^{*})^{\perp}$ (see Lemma 2.2). [^5] Multiplying (2.16) equation by $W$ yields that $\int WG=0$, which suggest that this is a solvability condition for $\Psi$. The equation (2.16) is therefore a suitable candidate for the Lyapunov-Schmidt reduction. We thus define the projection $Q\colon L^{2}\to L^{2}$ onto the range of $\frac{\mathrm{d}^{2}}{\mathrm{d}r^{2}}-(K_{0}+K_{2}\xi^{2})$:
$QF:=F-\left(\int FW\right)W^{*}.$ (2.17)
Indeed, if $\Upsilon^{\prime\prime}-(K_{0}+K_{2}\xi^{2})\Upsilon=F$ then multiplication by $W$ and integration by parts shows that $\int FW=0$. We can thus rewrite (2.15) as the projected equation